Need some help -

PCl5 can be produced by the reaction of phosphorus trichloride with excess chlorine as follows: PCl3 + Cl2 → PCl5

If the actual yield is 127 g of PCl5 and the percent yield is 84.8%, what mass of PCl3 must be used? Hint: Use actual yield and percent yield to find theoretical yield and proceed from there.


a. 150 g

b. 98.8 g

c.108 g

d.83.8 g

e.71.1 g

2. If the reaction N2 + 3 H2 → 2 NH3 is carried out using 1.40 g of N2 and 0.400 g of H2, what mass of excess reactant will remain?

a. 0.200 g N2

b. 1.00 g N2

c. Both reactants will be completely consumed

d. 0.100 g H2

e.0.300 g H2

3. How many moles of bromine will react with 0.0500 mole of C2H2 in the reaction C2H4 + Br2 → C2H4Br2?

a. 0.186 mol

b. 2.00 mol

c. 0.0250 mol

d. 0.100 mol

e. 0.0500 mol
Answer #1

Percentage yield =  actual yield Thcoreicaluied100 x 100 Theoreticalyield

  Theoretical yield-actualyield percentageyield × 100

1279 × 100 Theoretical yield

Theoretical yield- 149.76q

According to the equation PCl3 + Cl2\rightarrow PCl5

One mole of PCl3 should produce one mole of PCl5 when theoretical yeild is calculated it is assumed that the number of moles PCl5 will be formed will be equal to the number of moles of the limiting reactant ( reactant present in lesser concentration )

Therefore number of moles of PCl5 formed theoretically = number of moles of PCl3

Molar mass of PCl5 = 208.24 g / mol

theoretically number of moles of PCl5 =  theoretical yieldo f PCl5 molarmasso PC

=  149.76 208.24

= 0.719 moles

Therefore number of moles of PCl3 = 0.719 moles

molar mass of PCl3 = 137.33 g / mol

mass of PCl3 =  ( number of moles of PCl3 ) \times ( molar mass of PCl3 )

= 0.719 moles \times 137.33 g / mol

= 98.74 g

= 98.8 g ( b )

2)

According to the equation N2 + 3 H2\rightarrow 2 NH3

one mole of N2 reacts with 3 moles of H2

mass of N2 = 1.40 g

molar mass of N2 = 14 g / mol

number of moles of N2 = massofN2 molarmasso f N2 \

= \frac{1.4 }{14 }

= 0.1 moles

mass of H2 = 0.40 g

molar mass of H2 = 2 g / mol

number of moles of H2 =  masso fH2 molarmasso f H2

= 4

= 0.2 moles

For N2 to completely react There should be 3 \times 0.1 = 0.3 moles of H2 since number of moles of H2  = 0.2 moles < 0.3 H2 will be the limiting reagent and N2 will be left

number of moles of N2 will be left = number of moles of N2 - numberofmoleso f H2

= 0.1 -02

= 0.1 - 0.067

= 0.033 moles

mass of N2 will be left = number of moles of N2 will be left \times molar mass of N2

= 0.033 \times 14

= 0.462 g

3)

C2H4 + Br2\rightarrow C2H4Br2

According to the equation one mole of C2H4 reacts with one mole of Br2

Therefore 0.0500 moles of C2H4 reacts with 0.0500 moles of Br2